# Arithmetic Properties of Generalized Cubic and Overcubic Partitions

### v1

@@ 2.2. Radu’s Algorithm
$\sum_{n\geq 0}\bar{a}_{2^{k}i+2^{k-1}-3}(4n+3)q^{n}$ $\equiv-2^{2}(2^{k-1}(6i+7)+2)\dfrac{\varphi\left(q^{2}\right)^{2^{k-1}(6i+7)+1}\psi\left(q^{2}\right)\psi\left(q^{4}\right)}{\varphi\left(-q^{2}\right)^{2^{k}(i+1)}}\pmod{2^{k+3}}.$
@@
$\sum_{n=0}^{\infty}\bar{a}_{9i}(n)q^{n}$ $=\dfrac{\varphi\left(-q^{9}\right)\varphi\left(-q^{18}\right)^{9i-1}}{\varphi\left(-q^{3}\right)^{4}\varphi\left(-q^{6}\right)^{36i-4}}\Bigg{(}\left(\varphi\left(-q^{9}\right)-2\omega q\Omega\left(-q^{3}\right)\right)\times\left(\varphi\left(-q^{9}\right)-2\omega^{2}q\Omega\left(-q^{3}\right)\right)$
×$\quad\times\left(\varphi\left(-q^{18}\right)-2\omega q^{2}\Omega\left(-q^{6}\right)\right)^{9i-1}\times\left(\varphi\left(-q^{18}\right)-2\omega^{2}q^{2}\Omega\left(-q^{6}\right)\right)^{9i-1}\Bigg{)}.$
@@
×$\quad\times\left(\varphi\left(-q^{18}\right)^{2}+2q^{2}\varphi\left(-q^{18}\right)\Omega\left(-q^{6}\right)\right)^{9i-1}\Bigg{)}\pmod{4}$
@@
×$\quad\times\left(\varphi\left(-q^{18}\right)^{18i-2}+2(9i-1)q^{2}\varphi\left(-q^{18}\right)^{18i-3}\Omega\left(-q^{6}\right)\right)\Bigg{)}\pmod{4}.$
@@
(6.1) $G(\tau):=\frac{\eta(\delta_{1}\tau)^{r_{1}}\eta(\delta_{2}\tau)^{r_{2}}\cdots\eta(\delta_{u}\tau)^{r_{u}}}{\eta(\gamma_{1}\tau)^{s_{1}}\eta(\gamma_{2}\tau)^{s_{2}}\cdots\eta(\gamma_{t}\tau)^{s_{t}}},$
where $r_{i},s_{i},\delta_{i},$ and $\gamma_{i}$ are positive integers with $\delta_{1},...,\delta_{u},\gamma_{1},...,\gamma_{t}$ distinct, $u,t\geq 0$. The weight of $G(\tau)$ is given by
@@
(6.2) $p^{a}\geq\sqrt{\frac{\sum_{i=1}^{t}\gamma_{i}s_{i}}{\sum_{i=1}^{u}\dfrac{r_{i}}{\delta_{i}}}},$
@@
Also $\mathcal{D_{G}}=4$ and the weight is $-c/2\in\mathbb{Z}$ iff $c=2^{\alpha}m$ with $\alpha\geq 1$ and $\gcd(2,m)=1$. Next, we see that $42^{2}|$ 4 and
$4\displaystyle 2^{2}\geq\sqrt{\frac{2+(4c-6)}{\dfrac{c-1}{4}}}=\sqrt{16}=$ 4 .
@@
The proofs of Theorems 1.14 and 1.15 do not follow from Theorem 6.1. We present them below.
@@
(6.3) $\sum_{n\geq 0}\bar{a}_{2^{\alpha}m}(n)q^{n}=\frac{f_{4}^{2^{\alpha}m-1}}{f_{1}^{2}f_{2}^{2^{\alpha+1}m-3}}.$
@@
(6.4) $A^{3^{k}}_{\alpha,m}(z):=\frac{\eta^{3^{k+1}}(24z)}{\eta^{3^{k}}(72z)}\equiv 1\pmod{3^{k+1}}.$
@@
(6.5) $B_{\alpha,m,k}(z)$ $=\frac{\eta^{2^{\alpha}m-1}(96z)}{\eta^{2}(24z)\eta^{2^{\alpha+1}m-3}(48z)}\equiv\frac{f_{96}^{2^{\alpha}m-1}}{f_{24}^{2}f_{48}^{2^{\alpha+1}m-3}}\equiv\sum_{n=0}^{\infty}\bar{a}_{2^{\alpha}m}(n)q^{24n}\pmod{3^{k+1}}.$
@@
$3M\cdot\left(4\left(3^{k}-3^{k-2}\right)+2^{\alpha}m-1\right)\equiv 0\pmod{24}.$
@@
$\left(3^{k+1}-2\right)\frac{\gcd(d,24)^{2}}{24}-\left(2^{\alpha+1}m-3\right)\frac{\gcd(d,48)^{2}}{48}-3^{k}\frac{\gcd(d,72)^{2}}{72}+\left(2^{\alpha}m-1\right)\frac{\gcd(d,96)^{2}}{96}\geq 0.$
@@
$8\cdot 3^{k}-9\cdot 2^{\alpha}m+9\geq 0$
@@
Proof of Theorem 1.15.
Consider $t=p_{1}^{a_{1}}\cdots p_{m}^{a_{m}}$ where $p_{i}$’s are prime numbers greater than or equal to $5$. Then, we have
(6.6) $\bar{a}_{t}(n)q^{n}=\frac{f_{4}^{t-1}}{f_{1}^{2}f_{2}^{2t-3}}.$
We define
$G_{i}(z):=\frac{\eta^{p_{i}^{a_{i}}}(24z)}{\eta(24p_{i}^{a_{i}}z)}.$
From the binomial theorem, for any integer $k\geq 1$, and prime $p$ we have
(6.7) $G^{p_{i}^{k}}_{i}(z):=\frac{\eta^{p_{i}^{a_{i}+k}}(24z)}{\eta^{p_{i}^{k}}(24p_{i}^{a_{i}}z)}\equiv 1\pmod{p_{i}^{k+1}}.$
Next we define,
$H_{i,k,t}(z)$ $=\frac{\eta^{t-1}(96z)}{\eta^{2}(24z)\eta^{2t-3}(48z)}A^{p_{i}^{k}}_{i}(z)=\frac{\eta^{p_{i}^{a_{i}+k}-2}(24z)\eta^{t-1}(96z)}{\eta^{2t-3}(48z)\eta^{p_{i}^{k}}(24p_{i}^{a_{i}}z)}.$
From (6.6) and (6.7), we have
(6.8) $H_{i,k,t}(z)=\frac{\eta^{t-1}(96z)}{\eta^{2}(24z)\eta^{2t-3}(48z)}\equiv\frac{f_{96}^{t-1}}{f_{24}^{2}f_{48}^{2t-3}}\equiv\sum_{n=0}^{\infty}\bar{a}_{t}(n)q^{24n}\pmod{p_{i}^{k+1}}.$
Now, we will prove that $H_{i,k,t}(z)$ is a modular form. From Theorem 2.2, we find that the level of $H_{i,k,t}(z)$ is $N=96M$, where $M$ is the smallest positive integer such that
$96M\left(\frac{p_{i}^{a_{i}+k}-2}{24}-\frac{2t-3}{48}+\frac{t-1}{96}-\frac{p_{i}^{k}}{24p_{i}^{a_{i}}}\right)\equiv 0\pmod{24}$
which implies
$M\cdot\left(4p_{i}^{k}\left(p_{i}^{a_{i}}-\frac{1}{p_{i}^{a_{i}}}\right)-3(t+1)\right)\equiv 0\pmod{24}.$
Therefore, $M=2^{3}3$ and $N=2^{8}3^{2}$. The cusps of $\Gamma_{0}(2^{8}3^{2})$ are given by fractions $c/d$ where $d|2^{8}3^{2}$ and $\gcd(c,d)=1$. By Theorem 2.2, $H_{i,k,t}(z)$ is holomorphic at a cusp $c/d$ if and only if
$\left(p_{i}^{a_{i}+k}-2\right)\frac{\gcd(d,24)^{2}}{24}-\left(2t-3\right)\frac{\gcd(d,48)^{2}}{48}+\left(t-1\right)\frac{\gcd(d,96)^{2}}{96})-p_{i}^{k}\frac{\gcd(d,24p_{i}^{a_{i}})^{2}}{24p_{i}^{a_{i}}}\geq 0.$
Equivalently, $H_{i,k,t}(z)$ is holomorphic at a cusp $c/d$ if and only if the following inequalities are satisfied for the different values of $d$ as mentioned in the table below (note that we have simplified the inequalities as far as possible).
Due to our hypothesis, all the above inequalities in the list are already satisfied.
Therefore, $H_{i,k,t}(z)$ is holomorphic at a cusp $c/d$. The weight of $H_{i,k,t}(z)$ is $\ell=\frac{1}{2}\left(p_{i}^{k}\left(p_{i}^{a_{i}}-1\right)-t\right)$, which is a positive integer, and the associated character is given by
$\chi(\bullet)=\left(\frac{(-1)^{\ell}2^{3p_{i}^{a_{i}}-3t+1}3^{p_{i}^{a_{i}}-t}\left(p_{i}^{a_{i}}\right)^{-p_{i}^{k}}}{\bullet}\right).$
Thus, $H_{i,k,t}(z)\in M_{\frac{1}{2}\left(p_{i}^{k}\left(p_{i}^{a_{i}}-1\right)-t\right)}\left(\Gamma_{0}(2^{8}3^{2}),\chi(\bullet)\right)$. Thus by Theorem 2.5, the Fourier coefficients of $H_{i,k,t}(z)$ are almost divisible by $r=p_{i}^{k}$. Due to (6.8), this holds for $\bar{a}_{t}$ also. This completes the proof. ∎
@@
(6.9) $B_{\alpha,m,k}:=\sum_{n=0}^{\infty}\mathcal{B}_{\alpha,m,k}(n)q^{n}\equiv\sum_{n=0}^{\infty}\bar{a}_{2^{\alpha}m}\left(\frac{n}{24}\right)q^{n}\pmod{3^{k+1}}.$
Now, $B_{\alpha,m,k}(z)\in M_{3^{k}-2^{\alpha-1}m}\left(\Gamma_{0}(2^{8}3),\chi(\bullet)\right)$, where $\chi(\bullet)$ is the associated character. From the Theorem 6.2, we get that there is an integer $i\geq 0$ such that for any $j\geq 1$,
@@
(6.10) $\mathcal{B}_{\alpha,m,k}(q_{1}\cdots q_{i+j}\cdot n)\equiv 0\pmod{3^{j}}$
@@ 7. Concluding Remarks
    (7.1) $a_{25}(31^{2}n+644)$ $\equiv 0\pmod{31^{2}},$
    (7.2) $a_{41}(47^{2}n+256)$ $\equiv 0\pmod{47^{2}},$
    (7.3) $a_{61}(67^{2}n+555)$ $\equiv 0\pmod{67^{2}}.$
@@
    It can be noted that the congruences in Theorem 1.8 are valid for any arbitrary subscript $c$, not necessarily for $2i$. We restrict to the given case for ease in notations while writing the proof.
  5. (5)
    Theorem 1.8 can be further strengthened to the following: for all $n\geq 0$ and $i\geq 1$, we have
    $\bar{a}_{2i}(8n+5)$ $\equiv 0\pmod{8},$
    $\bar{a}_{2i}(8n+7)$ $\equiv 0\pmod{16}.$
    We leave the relevant details for the reader.
  6. (6)

### v2

@@ 2.2. Radu’s Algorithm
$\sum_{n\geq 0}\bar{a}_{2^{k}i+2^{k-1}-3}(4n+3)q^{n}$ $\equiv-2^{2}(2^{k-1}(6i+7)+2)$
$\quad\times\dfrac{\varphi\left(q^{2}\right)^{2^{k-1}(6i+7)+1}\psi\left(q^{2}\right)\psi\left(q^{4}\right)}{\varphi\left(-q^{2}\right)^{2^{k}(i+1)}}\pmod{2^{k+3}}$.
@@
$\sum_{n=0}^{\infty}\bar{a}_{9i}(n)q^{n}$ $=\dfrac{\varphi\left(-q^{9}\right)\varphi\left(-q^{18}\right)^{9i-1}}{\varphi\left(-q^{3}\right)^{4}\varphi\left(-q^{6}\right)^{36i-4}}\Bigg(\left(\varphi\left(-q^{9}\right)-2\omega q\Omega\left(-q^{3}\right)\right)$
$\quad\times\left(\varphi\left(-q^{9}\right)-2\omega^{2}q\Omega\left(-q^{3}\right)\right)\left(\varphi\left(-q^{18}\right)-2\omega q^{2}\Omega\left(-q^{6}\right)\right)^{9i-1}$
$\quad\times\left(\varphi\left(-q^{18}\right)-2\omega^{2}q^{2}\Omega\left(-q^{6}\right)\right)^{9i-1}\Bigg).$
@@
$\quad\times\left(\varphi\left(-q^{18}\right)^{2}+2q^{2}\varphi\left(-q^{18}\right)\Omega\left(-q^{6}\right)\right)^{9i-1}\Bigg)\pmod{4}$
@@
$\quad\times\left(\varphi\left(-q^{18}\right)^{18i-2}+2(9i-1)q^{2}\varphi\left(-q^{18}\right)^{18i-3}\Omega\left(-q^{6}\right)\right)\Bigg)\pmod{4}$.
@@
$G(\tau):=\frac{\eta(\delta_{1}\tau)^{r_{1}}\eta(\delta_{2}\tau)^{r_{2}}\cdots\eta(\delta_{u}\tau)^{r_{u}}}{\eta(\gamma_{1}\tau)^{s_{1}}\eta(\gamma_{2}\tau)^{s_{2}}\cdots\eta(\gamma_{t}\tau)^{s_{t}}}=q^{\tfrac{E_{G}}{24}}\sum\limits_{n=0}^{\infty}b(n)q^{n},$ (6.1)
where $r_{i},s_{i},\delta_{i},$ and $\gamma_{i}$ are positive integers with $\delta_{1},...,\delta_{u},\gamma_{1},...,\gamma_{t}$ distinct, $u,t\geq 0$, and
$E_{G}:=\sum\limits_{i=1}^{u}\delta_{i}r_{i}-\sum\limits_{i=1}^{t}\gamma_{i}s_{i}.$
The weight of $G(\tau)$ is given by
@@
$p^{a}\geq\sqrt{\frac{\sum_{i=1}^{t}\gamma_{i}s_{i}}{\sum_{i=1}^{u}\dfrac{r_{i}}{\delta_{i}}}},$ (6.2)
@@
Also $\mathcal{D_{G}}=4$ and the weight is $c/2\in\mathbb{Z}$ iff $c=2^{\alpha}m$ with $\alpha\geq 1$ and $\gcd(2,m)=1$. Next, we see that $2^{2}|4$ and
$2^{2}\geq\sqrt{\frac{2+(4c-6)}{\dfrac{c-1}{4}}}=\sqrt{16}=4.$
@@
The proof of Theorem 1.10 do not follow from Theorem 6.1. We present it below.
@@
$\sum_{n\geq 0}\bar{a}_{2^{\alpha}m}(n)q^{n}=\frac{f_{4}^{2^{\alpha}m-1}}{f_{1}^{2}f_{2}^{2^{\alpha+1}m-3}}.$ (6.3)
@@
$A^{3^{k}}_{\alpha,m}(z):=\frac{\eta^{3^{k+1}}(24z)}{\eta^{3^{k}}(72z)}\equiv 1\pmod{3^{k+1}}$. (6.4)
@@
$B_{\alpha,m,k}(z)$ $=\frac{\eta^{2^{\alpha}m-1}(96z)}{\eta^{2}(24z)\eta^{2^{\alpha+1}m-3}(48z)}=\frac{f_{96}^{2^{\alpha}m-1}}{f_{24}^{2}f_{48}^{2^{\alpha+1}m-3}}\equiv\sum_{n=0}^{\infty}\bar{a}_{2^{\alpha}m}(n)q^{24n}\pmod{3^{k+1}}$. (6.5)
@@
$3M\cdot\left(4\left(3^{k}-3^{k-2}\right)-2^{\alpha}m-1\right)\equiv 0\pmod{24}$.
@@
$\left(3^{k+1}-2\right)\frac{\gcd(d,24)^{2}}{24}-\left(2^{\alpha+1}m-3\right)\frac{\gcd(d,48)^{2}}{48}-3^{k}\frac{\gcd(d,72)^{2}}{72}\\ +\left(2^{\alpha}m-1\right)\frac{\gcd(d,96)^{2}}{96}\geq 0.$
@@
Value(s) of $d$ Inequality
$1,2,4,8,24$ $32\cdot 3^{k}-9\cdot 2^{\alpha}m-9\geq 0$
$3,6,12$ $3^{k}-\dfrac{9}{32}(2^{\alpha}m+1)\geq 0$
$16,48$ $8\cdot 3^{k}-9\cdot 2^{\alpha}m+9\geq 0$
$32,64,128,256$ $3^{k-2}\geq 0$
$96,192,384,768$ $3^{k}\geq 0$
@@
$8\cdot 3^{k}-9\cdot 2^{\alpha}m+9\geq 0.$
@@
$B_{\alpha,m,k}:=\sum_{n=0}^{\infty}\mathcal{B}_{\alpha,m,k}(n)q^{n}\equiv\sum_{n=0}^{\infty}\bar{a}_{2^{\alpha}m}\left(\frac{n}{24}\right)q^{n}\pmod{3^{k+1}}$. (6.6)
Now, $B_{\alpha,m,k}(z)\in M_{3^{k}-2^{\alpha-1}m}\left(\Gamma_{0}(2^{8}3),\chi(\bullet)\right)$, where $\chi(\bullet)$ is the associated character. From Theorem 6.2, we get that there is an integer $i\geq 0$ such that for any $j\geq 1$,
@@
$\mathcal{B}_{\alpha,m,k}(q_{1}\cdots q_{i+j}\cdot n)\equiv 0\pmod{3^{j}}$ (6.7)
@@ 7. Concluding Remarks
    $a_{25}(31^{2}n+644)$ $\equiv 0\pmod{31^{2}}$, (7.1)
    $a_{41}(47^{2}n+256)$ $\equiv 0\pmod{47^{2}}$, (7.2)
    $a_{61}(67^{2}n+555)$ $\equiv 0\pmod{67^{2}}$. (7.3)
@@
    Congruences $(m,M,N,t,(r_{\delta}))$ and $(r_{\delta}^{\prime})$ $P(t)$ $\lfloor\nu\rfloor$
    (7.1) (961,62,62,644,(960,24,31,0))(961,62,62,644,(960,-24,-31,0)) $\{644\}$ $3813$
    and $(49,0,0,0)$
    (7.2) (2209,94,94,256,(2208,40,47,0))(2209,94,94,256,(2208,-40,-47,0)) $\{256\}$ $13208$
    and $(81,0,0,0)$
    (7.3) (4489,134,134,555,(4488,60,67,0))(4489,134,134,555,(4488,-60,-67,0)) $\{555\}$ $38091$
    and $(121,0,0,0)$
@@
    After this paper was posted on ArXiv, Ghoshal and Jana [GJ25] announced a proof of this conjecture.
Declarations
The authors make the following declaration.
  1. (1)
  2. (2)
    The authors have no relevant financial or non-financial interests to disclose.
  3. (3)
    There is no data associated with this manuscript.
  4. (4)
    All authors contributed to the study conception and design. All authors read and approved the final manuscript.